Determine the domain of the function f(x)=-x-10/-3x^3+5x^2+10x


1
Expert's answer
2022-01-06T15:05:55-0500
"f(x)=\dfrac{-x-10}{-3x^3+5x^2+10x}"

"-3x^3+5x^2+10x\not=0"

"-x(3x^2-5x-10)\not=0"

Let "3x^2-5x-10=0"


"D=(-5)^2-4(3)(-10)=145"

"x=\dfrac{5\pm\sqrt{145}}{2(3)}"

"x_1=\dfrac{5-\sqrt{145}}{6}, x_2=\dfrac{5+\sqrt{145}}{6}"

Domain:


"(-\infin, \dfrac{5-\sqrt{145}}{6})\cup(\dfrac{5-\sqrt{145}}{6}, 0)"

"\cup (0, \dfrac{5+\sqrt{145}}{6})\cup(\dfrac{5+\sqrt{145}}{6}, \infin)"